OpenStudy (prince1342):

Reggie can line a football field in 120 minutes. Rosalinda can line a football field in 80 minutes. If they work together, how many minutes does it take them to line a football field?

OpenStudy (prince1342):

40 minutes 48 minutes 200 minutes 240 minutes

OpenStudy (luffingsails):

This is sort of a unitless question. Assume that they need to line 100 yards. So, Reggie's rate is 100 yards/120 minutes = 5yards/6 minutes Rosalinda's rate is 100 yards/80 minutes = 5 yards / 4 minutes If they work together to line 100 yards, how long does it take them. 5/6 (X) + 5/4(X) = 100 yards 10/12(X) + 15/12(X) = 100 yards 25/12(X) = 100 yards 25(X) = 1200 (X) = 48 So, the time is 48 minutes.

OpenStudy (prince1342):

thank you so much

OpenStudy (luffingsails):

Welcome!